### v2

# FC-KAN: Function Combinations in Kolmogorov-Arnold Networks

Abstract
In this paper, we introduce FC-KAN, a Kolmogorov-Arnold Network (KAN) that leverages combinations of popular mathematical functions such as B-splines, wavelets, and radial basis functions on low-dimensional data through element-wise operations. We explore several methods for combining the outputs of these functions, including sum, element-wise product, the addition of sum and element-wise product, representations of quadratic and cubic functions, concatenation, linear transformation of the concatenated output, and others. In our experiments, we compare FC-KAN with a multi-layer perceptron network (MLP) and other existing KANs, such as BSRBF-KAN, EfficientKAN, FastKAN, and FasterKAN, on the MNIST and Fashion-MNIST datasets. A variant of FC-KAN, which uses a combination of outputs from B-splines and Difference of Gaussians (DoG) in the form of a quadratic function, outperformed all other models on the average of 5 independent training runs. We expect that FC-KAN can leverage function combinations to design future KANs. Our repository is publicly available at: https://github.com/hoangthangta/FC_KAN.
Keywords Kolmogorov Arnold Networks  $\cdot$ function combinations  $\cdot$ B-splines  $\cdot$ wavelets  $\cdot$ radial basis functions
1 Introduction
The work by Liu et al. [1, 2] helped to revive the usage of the Kolmogorov-Arnold representation theorem (KART) in neural networks. They highlighted the use of learnable activation functions as "edges" to fit training data better, in contrast to the traditional use of fixed activation functions as "nodes" in multi-layer perceptrons (MLPs). The foundation of KANs is based on KART, which was developed to address Hilbert’s 13th problem [3], specifically, the statement, “Prove that the equation of seventh degree $x^{7}+ax^{3}+bx^{2}+cx+1=0$ is not solvable by means of any continuous functions of only two variables”. KART asserts that any continuous function of multiple variables can be represented as a sum of continuous functions of a single variable [4]. Motivated by this idea, numerous researchers have worked on creating different types of KANs utilizing widely known polynomials and basis functions.
While these works use a single function (not mentioning activation functions) in constructing KANs [5, 6, 7, 8], there are a few ones using function combinations [9, 10, 11, 2]. In several works that designed MLP-based neural networks, the function combinations typically occur in activation functions to improve model performance and stability [12, 13, 14]. In KANs, the combinations occur directly on the functions used to fit input data, rather than through activation functions. For instance, BSRBF-KAN [9], which combines B-splines and Gaussian Radial Basis Functions (GRBFs) in all network layers. However, the authors only applied addition operations to data tensors without incorporating element-wise multiplications, which we believe does not effectively capture data features. Additionally, element-wise multiplications may lead to memory issues or slow calculations when performed on high-dimensional data. While Liu et al. [2] used element-wise multiplications alongside additions in MultKAN to enhance data capture, their focus was limited to small-scale examples.
With the purpose of exploiting data features efficiently, we propose a novel KAN, FC-KAN (Function Combinations in Kolmogorov-Arnold Networks), which leverages various functions to capture input data throughout the network layers and combines them in low-dimensional spaces, such as the output layer, using various methods mainly based on element-wise operations, including sum, product, the combination of sum and product, representations of quadratic and cubic functions, and concatenation. We avoid using a higher-degree function due to their increased computational demands on data tensors, which can lead to memory errors. As a result, the model is able to capture more data features, leading to improved performance on the MNIST and Fashion-MNIST datasets compared to other KAN networks. Additionally, employing n-degree functions aligns with the core concept of KAN, where they are used both to capture input data and to represent data features in the output. We expect this exciting development to drive the proliferation of basis function combinations in neural networks. In summary, our main contributions are:
  • Construct FC-KAN to demonstrate the effectiveness of function combinations through various methods applied to low-dimensional data, such as model outputs.
  • Compare combination methods on the function outputs using mainly element-wise operations to search for proper methods.
Aside from this section, the paper is organized as follows: Section 2 discusses related work on KART and KANs. Section 3 details our methodology, covering KART, the design of the KAN architecture, several existing KANs, and FC-KAN. Section 4 presents our experiments, comparing FC-KAN variants with MLP and other existing KANs using data from the MNIST and Fashion-MNIST datasets. Additionally, this section includes a comparison of combination methods within FC-KAN. We address some limitations of our study in Section 5. Lastly, Section 6 offers our conclusions and potential directions for future research.
2 Related Works
2.1 KART and KAN
In 1957, Kolmogorov provided a proof to Hilbert’s 13th problem by showing that any multivariate continuous function can be expressed as a combination of single-variable functions and additions, a concept known as KART [4, 15]. This theorem has been utilized in many studies to develop neural networks [16, 17, 18, 19, 20, 21]. However, there is an ongoing debate about the applicability of KART in neural network design. Girosi and Poggio [22] argued that KART’s relevance to neural networks is questionable because the inner function $\phi_{q,p}$ in Equation 1 may be highly non-smooth [23], which could hinder $f$ from being smooth—a key attribute for generalization and noise resistance in neural networks. Conversely, Kůrková [24] contended that KART is applicable to neural networks, showing that linear combinations of affine functions can effectively approximate all single-variable functions using certain sigmoidal functions.
Despite the long history of KART’s application in neural networks, it had not garnered significant attention in the research community until the recent study by Liu et al. [1]. They suggested moving away from strict adherence to KART and generalizing it to develop KANs with additional neurons and layers. Our intuition aligns with this perspective as it helps to mitigate the issue of non-smooth functions when applying KART to neural networks. Consequently, KANs have the potential to outperform MLPs in both accuracy and interpretability for small-scale AI + Science tasks. However, the resurgence of KANs has also faced criticism from Dhiman [25], who argue that KANs are essentially MLPs with spline-based activation functions, in contrast to traditional MLPs with fixed activation functions.
By introducing a new perspective to the scientific community in the new design of neural networks, KANs inspired many works to prove their effectiveness by topics, including expensive problems [26], keyword spotting [27], mechanics problems [28], quantum computing [29, 30, 31], survival analysis [32], time series forecasting [33, 34, 35, 36, 37], and vision tasks [38, 39, 40]. Also, many novel KANs utilize well-known mathematical functions, particularly those capable of handling curves, such as B-Splines [41] (Original KAN [1], EfficientKAN [^1] https://github.com/Blealtan/efficient-kan, BSRBF-KAN [9]), Gaussian Radial Basis Functions (GRBFs) (FastKAN [5], DeepOKAN [28], BSRBF-KAN [9]), Reflection SWitch Activation Function (RSWAF) in FasterKAN [6], Chebyshev polynomials (TorchKAN [42], Chebyshev KAN [8]), Legendre polynomials (TorchKAN [42]), Fourier transform (FourierKAN [^2] https://github.com/GistNoesis/FourierKAN/, FourierKAN-GCF [43]), wavelets [7, 44], and other polynomial functions [45].
2.2 Function combinations in KANs and other neural networks
Several works utilize the function combinations to design novel KANs. Ta [9] mentioned the combination of functions – B-Splines and radial basis functions – in designing KANs. Their BSRBF-KAN showed better convergence on the training data for MNIST and Fashion-MNIST. Liu et al. [2] introduced MultKAN, which consists of multiplication operations to capture multiplicative structures in data better. Unlike KAN [1], which directly copies nodes, MultKAN uses both addition nodes (copied from subnodes) and multiplication nodes (multiplying multiple subnodes). However, their work focused on small-scale examples only. Yang et al. [10] utilized function combinations to create optimal activation functions at each node using an adaptive strategy, addressing the drawbacks of single activation functions in their S-KAN model. They also extended S-KAN to S-ConvKAN, which showed superior performance in image classification tasks, outperforming CNNs and KANs with comparable structures. In another work,  Altarabichi [11] proposed the replacement of the sum with the average function in KAN neurons that can improve the model performance and keep the training stability in their DropKAN [46].
Before the appearance of KANs, several studies focused on constructing combinations of activation functions (along with their weights) or utilizing a set of flexible activations in neural networks to enhance model performance and stability. Jie et al. [12] introduced a new family of flexible activation functions for LSTM cells, along with another family developed by combining ReLUs and ELUs. Their findings demonstrate that LSTM models using the P-Sig-Ramp flexible activations significantly improve time series forecasting. Additionally, the P-E2-ReLU activation exhibits enhanced stability and performance in lossy image compression tasks with convolutional autoencoders. Xu and Chen [13] investigated a selection of widely used activation functions across various datasets in classification and regression tasks. They then created combinations of the best-performing activation functions within a convex restriction, showing improved performance compared to the corresponding base activation functions in a standard broad learning system. In another study, Zhang [14] developed a novel deep neural network (DNN) that optimizes activation function combinations for different neurons based on extensive simulations. The experiments showed that their DNNs outperformed those that rely on a single activation function. Instead of using the inner product between data vectors and weights,  Fan et al. [47] replaced this operation with a quadratic function in neurons within deep quadratic neural networks. Quadratic neurons provide enhanced expressive capability compared to conventional neurons, highlighting the advantages of quadratic networks in terms of expressive efficiency, unique representation, compact architecture, and computational capacity.
In addition to exploring function combinations in mathematics, it is also possible to merge function outputs using tensor operations. When the data dimensions are the same, element-wise operations such as summation or product can be applied to combine these outputs [48] which appear in many multi-modal tasks [49, 50, 51]. However, the computational challenges posed by high-dimensional data, such as inefficiencies in tensor-product operations, have led to a growing body of work aimed at optimizing and accelerating these processes [52, 53]. Furthermore, other studies have focused on tensor fusion and tensor decomposition techniques to extract meaningful features from data tensors and to simplify data combinations [54, 55, 56, 57].
3 Methodology
3.1 Kolmogorov-Arnold Representation Theorem
A KAN is based on KART, which states that any continuous multivariate function $f$ defined on a bounded domain can be represented as a finite combination of continuous single-variable functions and their additions [58, 59]. For a set of variables $\mathbf{x}={x_{1},x_{2},\ldots,x_{n}}$, where $n$ is the number of variables, the multivariate continuous function $f(\mathbf{x})$ is expressed as:
$f(\mathbf{x})=f(x_{1},\ldots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right)$ (1)
which has two types of summations: the outer sum and the inner sum. The outer sum, $\sum_{q=1}^{2n+1}$, aggregates $12n+$ 1 terms of $\Phi_{q}$ ($\mathbb{R}\to\mathbb{R}$). The inner sum, on the other hand, aggregates $n$ terms for each $q$, where each term $\phi_{q,p}$ ($\phi_{q,p}\colon[0,1]\to\mathbb{R}$) denotes a continuous function of a single variable $x_{p}$.
3.2 The design of KANs
Remind an MLP that consists of affine transformations and non-linear functions. Starting with an input $\mathbf{x}$, the network processes it through a series of weight matrices across layers (from layer 00 to layer $L-1$) and applies the non-linear activation function $\sigma$ to produce the final output.
$\text{MLP}(\mathbf{x})$ $=(W_{L-1}\circ\sigma\circ W_{L-2}\circ\sigma\circ\cdots\circ W_{1}\circ\sigma\circ W_{0})\mathbf{x}$ (2)
$=\sigma\left(W_{L-1}\sigma\left(W_{L-2}\sigma\left(\cdots\sigma\left(W_{1}\sigma\left(W_{0}\mathbf{x}\right)\right)\right)\right)\right)$
Inspired by KART, Liu et al. [1] developed KANs but recommended extending the approach to incorporate greater network widths and depths. To address this, appropriate functions $\Phi_{q}$ and $\phi_{q,p}$ need to be identified. A typical KAN network with $L$ layers processes the input $\mathbf{x}$ to produce the output as follows:
$\text{KAN}(\mathbf{x})=(\Phi_{L-1}\circ\Phi_{L-2}\circ\cdots\circ\Phi_{1}\circ\Phi_{0})\mathbf{x}$ (3)
which $\Phi_{l}$ is the function matrix of the $l^{th}$ KAN layer or a set of pre-activations. Let denote the neuron $i^{th}$ of the layer $l^{th}$ and the neuron $j^{th}$ of the layer $l+1^{th}$. The activation function $\phi_{l,i,j}$ connects $(l,i)$ to $(l+1,j)$:
$\phi_{l,j,i},\quad l=0,\cdots,L-1,\quad i=1,\cdots,n_{l},\quad j=1,\cdots,n_{l+1}$ (4)
with $n_{l}$ is the number of nodes of the layer $l^{th}$. So now, the function matrix $\Phi_{l}$ can be represented as a matrix $n_{l+1}\times n_{l}$ of activations as:
$\mathbf{x}_{l+1}=\underbrace{\left(\begin{array}[]{cccc}\phi_{l,1,1}(\cdot)&\phi_{l,1,2}(\cdot)&\cdots&\phi_{l,1,n_{l}}(\cdot)\\ \phi_{l,2,1}(\cdot)&\phi_{l,2,2}(\cdot)&\cdots&\phi_{l,2,n_{l}}(\cdot)\\ \vdots&\vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\cdot)&\phi_{l,n_{l+1},2}(\cdot)&\cdots&\phi_{l,n_{l+1},n_{l}}(\cdot)\end{array}\right)}_{\Phi_{l}}\mathbf{x}_{l}$ (5)
3.3 Implementation of the current KANs
Figure 1: Left: The structure of KAN(2,3,1). Right: The simulation of how to calculate $\phi_{1,1,1}$ by control points and B-splines.
LiuKAN [^3] We refer to the original KAN as LiuKAN, following the first author’s last name [1], while another work [7] refers to it as Spl-KAN. was implemented by Liu et al. [1] by using the residual activation function $\phi(x)$ as the sum of the base function and the spline function with their corresponding weight matrices $w_{b}$ and $w_{s}$:
$\phi(x)=w_{b}b(x)+w_{s}spline(x)$ (6)
$b(x)=silu(x)=\frac{x}{1+e^{-x}}$ (7)
$spline(x)=\sum_{i}c_{i}B_{i}(x)$ (8)
where $b(x)$ equals to $silu(x)$ as in Equation 7 and $spline(x)$ is expressed as a linear combination of B-splines $B_{i}$s and control points (coefficients) $c_{i}$s as in Equation 8. Each activation function is activated with $w_{s}=1$ and $spline(x)\approx 0$, while $w_{b}$ is initialized by using Xavier initialization. Figure 1 shows the structure of KAN(2,3,1) and how to calculate an inner function $\phi$ by control points and B-splines.
EfficientKAN adopted the same approach as Liu et al. [1] but reworked the computation using B-splines followed by linear combination, reducing memory cost and simplifying computation [^4] https://github.com/Blealtan/efficient-kan. The authors replaced the incompatible L1 regularization on input samples with L1 regularization on weights. They also added learnable scales for activation functions and switched the base weight and spline scaler matrices to Kaiming uniform initialization, significantly improving performance on MNIST.
FastKAN can speed up training over EfficientKAN by using GRBFs to approximate the 3-order B-spline and employing layer normalization to keep inputs within the RBFs’ domain [5]. These modifications simplify the implementation without sacrificing accuracy. The RBF has the formula:
$\phi(r)=e^{-\epsilon r^{2}}$ (9)
where $r=\|x-c\|$ is the distance between an input vector $x$ and a center $c$, and $\epsilon$ ($epsilon>0$) is a sharp parameter that controls the width of the Gaussian function. FastKAN uses a special form of RBFs, Gaussian RBFs where $\epsilon=\frac{1}{2}$ as [5]:
[…]
and $h$ for controlling the width of the Gaussian function. Finally, the RBF network with $N$ centers can be shown as [5]:
[…]
where $w_{i}$ represents adjustable weights or coefficients, and $\phi$ denotes the radial basis function as in Equation 9.
FasterKAN outperforms FastKAN in both forward and backward processing speeds [6]. It uses Reflectional Switch Activation Functions (RSWAFs), which are variants of RBFs and straightforward to compute due to their uniform grid structure. The RSWAF function is shown as follows:
[…]
Then, the RSWAF network with $N$ centers will be:
[…]
BSRBF-KAN is a KAN that combines B-splines from EfficientKAN and Gaussian RBFs from FastKAN in each network layer by additions [9]. It has a speedy convergence compared to EfficientKAN, FastKAN, and FasterKAN in training data. The BSRBF function is represented as:
$\phi_{BSRBF}(x)=w_{b}b(x)+w_{s}(\phi_{BS}(x)+\phi_{RBF}(x))$ (14)
where $b(x)$ and $w_{b}$ are the base function (linear) and its base matrix. $\phi_{BS}(x)$ and $\phi_{RBF}(x)$ are B-Spline and RBF functions, and $w_{s}$ is the spline matrix.
Wav-KAN is a neural network architecture that integrates wavelet functions into Kolmogorov-Arnold Networks to address challenges in interpretability, training speed, robustness, and computational efficiency found in MLP and LiuKAN [7]. By efficiently capturing both high and low-frequency components of input data, Wav-KAN achieves a balance between accurately representing the data structure and avoiding overfitting. The authors use several wavelet types, including the DoG, Mexican hat, Morlet, and Shannon. In our paper, we use the DoG function to combine other functions to create function combinations. The formula for DoG is:
$\psi(x)=\phi_{DOG}(x)=-\frac{d}{dx}\left(e^{-\frac{x^{2}}{2}}\right)=x\cdot e^{-\frac{x^{2}}{2}}$ (15)
which $\frac{d}{dx}$ is used to represent the derivative with respect to $x$. The term inside the derivative, $e^{-\frac{x^{2}}{2}}$, is a Gaussian function centered at 0.
3.4 FC-KAN
Ta [9] introduced the idea of combining functions, such as B-splines and GRBFs in BSRBF-KAN, to improve convergence when training models for image classification. However, their method was limited to element-wise addition of function outputs in each layer, without exploring other matrix operations like multiplications or different combinations. We argue that this approach might not effectively capture the input data’s features. It is important to note that multiplying matrices of high-dimensional data can lead to memory errors or inefficient running time on GPU/CPU devices. Therefore, it is wise to perform these operations on low-dimensional data, such as the output layer of a neural network, in data classification problems.
Figure 2: The structure of FC-KAN and the three types of combined outputs: element-wise, concatenation, and linearization.
We propose a novel network, FC-KAN (Function Combinations in Kolmogorov-Arnold Networks), which leverages function combinations applied to training data, considering the outputs as low-dimensional data. Given an input $\mathbf{x}$ and a set of functions $F=\{f_{1},f_{2},\dots,f_{n}\}$, where $n$ is the number of functions used, the input $\mathbf{x}$ is passed independently to each function $f_{i}$ through network layers, producing the output $\mathbf{o_{i}}$ as:
$\mathbf{o_{i}}=f_{i}(\mathbf{x})=(f_{i,L-1}\circ f_{i,L-2}\circ\cdots\circ f_{i,1}\circ f_{i,0})\mathbf{x}$ (16)
which $f_{i,l}$ is the function $f_{i}$ at the layer $l$. So we have a set of outputs $O=\{\mathbf{o_{1}},\mathbf{o_{2}},...,\mathbf{o_{n}}\}$ corresponding to the number of functions used. Note that all outputs in $O$ have the same size.
After that, we combine function outputs to obtain the element-wise output, the concatenated output, and the linearized output, as shown in Figure 2. The element-wise output is formed by using element-wise operations, including the sum output (Equation 17a), the element-wise product output (Equation 17b), the sum and element-wise product output (Equation 17c), the quadratic and cubic function outputs (Equation 17d and Equation 17e). The concatenated output is formed by simply concatenating the function outputs together as in Equation 17f. The linearized output is formed by passing the concatenated output through a linear transformation (with matrix multiplication by $W$ and addition of bias $b$) as in Equation 17g. The element-wise output and the linearized output will maintain the same data dimension as each element $\mathbf{o_{i}}$ in $O$, except in the concatenated output, where the data size will be $\mathbf{o_{i}}$ multiplied by the number of function outputs being combined.
$\mathbf{o}_{sum}=\sum_{i=1}^{n}\mathbf{o}_{i}=\mathbf{o}_{1}+\mathbf{o}_{2}+\cdots+\mathbf{o}_{n}$ (17a)
$\mathbf{o}_{prod}=\bigodot_{i=1}^{n}\mathbf{o}_{i}=\mathbf{o}_{1}\odot\mathbf{o}_{2}\odot\cdots\odot\mathbf{o}_{n}$ (17b)
$\mathbf{o}_{sum+prod}=\mathbf{o}_{sum}+\mathbf{o}_{prod}=\sum_{i=1}^{n}\mathbf{o}_{i}+\bigodot_{i=1}^{n}\mathbf{o}_{i}$ (17c)
$\begin{split}\mathbf{o}_{quad}=&\mathbf{o}_{sum+prod}+\sum_{i=1}^{n}\mathbf{o_{i}}\odot\mathbf{o_{i}}\\ =&\mathbf{o}_{sum+prod}+\mathbf{o}_{1}\odot\mathbf{o}_{1}+\mathbf{o}_{2}\odot\mathbf{o}_{2}+\dots+\mathbf{o}_{n}\odot\mathbf{o}_{n}\end{split}$ (17d)
$\begin{split}\mathbf{o}_{cubic}=\mathbf{o}_{quad}\odot\mathbf{o}_{sum}\end{split}$ (17e)
$\mathbf{o}_{concat}=concat(\mathbf{o}_{1},\mathbf{o}_{2},\cdots,\mathbf{o}_{n})$ (17f)
$\mathbf{o}_{concat\_linear}=W\cdot\mathbf{o}_{concat}+b$ (17g)
Figure 3: Various data combinations are performed using element-wise operations (additions $+$ and multiplications $\odot$) over two given outputs. The outputs always have the same data dimensions as the inputs.
Figure 3 presents data combinations over two given function outputs $\mathbf{o_{1}}$ and $\mathbf{o_{2}}$ using element-wise operations. The results are 5 combined outputs: sum, sum + prod, prod, quad, and cubic. Output combinations can utilize higher-degree functions, but these may significantly increase computational complexity, especially in matrix multiplication. Additionally, using more functions results in a larger number of outputs, which can further complicate data combination calculations. To manage this complexity, we prefer to restrict output combinations to quadratic functions involving up to two outputs. For instance, to combine DoG and B-Splines at the output, we can use the following quadratic function formula:
$\mathbf{o}_{DoG+BS}$ $=f_{DoG}(\mathbf{x})+f_{BS}(\mathbf{x})+f_{DoG}(\mathbf{x})\odot f_{BS}(\mathbf{x})+(f_{DoG}(\mathbf{x}))^{2}+(f_{BS}(\mathbf{x}))^{2}$ (18)
$=\cdots+\cdots+\cdots+f_{DoG}(\mathbf{x})\odot f_{DoG}(\mathbf{x})+f_{BS}(\mathbf{x})\odot f_{BS}(\mathbf{x})$
which $f_{DoG}$ and $f_{BS}$ refer to DoG and B-spline functions. Finally, we use the combined output to compute the cross-entropy loss against the true labels and train the models.
4 Experiments
4.1 Training Configuration
There are 5 independent training runs for each model on the MNIST [60] and Fashion-MNIST [61] datasets to obtain a more reliable overall performance assessment. We then select the best metric value from all runs to minimize the impact of training variability and accurately gauge the models’ maximum potential. To maintain simplicity in the network design, we utilized only activation functions (SiLU), linear transformations, and layer normalization in all models: BSRBF-KAN, EfficientKAN, FastKAN, FasterKAN, FC-KAN, and MLP. We do not use LiuKAN because its design, as the author intended, results in longer training times [1].
Table 1: The number of parameters by models.
Dataset Model Network structure #Params
MNIST + Fashion-MNIST BSRBF-KAN (784, 64, 10) 459040
FastKAN (784, 64, 10) 459114
FasterKAN (784, 64, 10) 408224
EfficientKAN (784, 64, 10) 508160
FC-KAN (784, 64, 10) 560820
MLP (784, 64, 10) 52512
Figure 4: The logarithmic values of training losses for the models over 25 epochs on MNIST and 35 epochs on Fashion-MNIST. A quadratic function is used to combine B-Splines and DoG at the output of FC-KAN.
As shown in Table 1, all models contain a network structure of (784, 64, 10), comprising 784 input neurons, 64 hidden neurons, and 10 output neurons corresponding to the 10 output classes (0-9). Due to the function combinations, FC-KAN has the highest number of parameters, while the MLP has the fewest because it only contains linear transformations over data between layers. The models were trained with 25 epochs on MNIST and 35 epochs on Fashion-MNIST. For KAN models, we use grid_size=5, spline_order=3, and num_grids=8. Other hyperparameters are the same in all models, including batch_size=64, learning_rate=1e-3, weight_decay=1e-4, gamma=0.8, optimize=AdamW, and loss=CrossEntropy.
In FC-KAN models, we combine 2 out of 4 functions: B-splines (denoted as BS), Radial Basis Functions (denoted as RBF, specifically using GRBFs), Difference of Gaussians (denoted as DoG), and base linear transformations (denoted as BASE), to create 6 FC-KAN variants. All variants use a quadratic function representation in the output for the experiments. The FC-KAN models are: FC-KAN (DoG+BS), FC-KAN (DoG+RBF), FC-KAN (DoG+BASE), FC-KAN (BS+RBF), FC-KAN (BS+BASE), and FC-KAN (RBF+BASE).
4.2 Model Performance
Figure 4 shows the training losses, represented on a logarithmic scale, for MLP and KAN models on the MNIST and Fashion-MNIST datasets. The loss performance of each model was evaluated based on an independent training run. FC-KAN consistently achieves the lowest training losses across both datasets, followed by BSRBF-KAN due to its fast convergence feature. In contrast, Faster-KAN records the highest training loss on MNIST, while MLP performs similarly on Fashion-MNIST.
Table 2: The average metric values in 5 training runs on MNIST and Fashion-MNIST. FC-KAN models use a quadratic function representation to combine outputs.
Dataset Model Train. Acc. Val. Acc. F1 Time (seconds)
MNIST BSRBF-KAN 100.00 ± 0.00 97.59 ± 0.02 97.56 ± 0.02 211.5
FastKAN 99.98 ± 0.01 97.47 ± 0.05 97.43 ± 0.05 164.47
FasterKAN 98.72 ± 0.02 97.69 ± 0.04 97.66 ± 0.04 161.88
EfficientKAN 99.40 ± 0.10 97.34 ± 0.05 97.30 ± 0.05 184.5
MLP 99.82 ± 0.08 97.74 ± 0.07 97.71 ± 0.07 146.58
FC-KAN (DoG+BS) 100.00 ± 0.00 97.91 ± 0.05 97.88 ± 0.05 263.29
FC-KAN (DoG+RBF) 100.00 ± 0.00 97.83 ± 0.04 97.81 ± 0.04 244.26
FC-KAN (DoG+BASE) 99.94 ± 0.04 97.81 ± 0.03 97.78 ± 0.03 234.13
FC-KAN (BS+RBF) 99.95 ± 0.03 97.74 ± 0.04 97.71 ± 0.04 234.07
FC-KAN (BS+BASE) 99.96 ± 0.02 97.78 ± 0.03 97.75 ± 0.03 235.04
FC-KAN (RBF+BASE) 99.97 ± 0.02 97.79 ± 0.03 97.76 ± 0.03 228.16
Fashion-MNIST BSRBF-KAN 99.34 ± 0.04 89.38 ± 0.06 89.36 ± 0.06 276.75
FastKAN 98.25 ± 0.07 89.40 ± 0.08 89.35 ± 0.08 208.68
FasterKAN 94.41 ± 0.03 89.31 ± 0.03 89.25 ± 0.02 220.7
EfficientKAN 94.81 ± 0.09 88.98 ± 0.07 88.91 ± 0.08 247.85
MLP 94.14 ± 0.04 88.94 ± 0.05 $8$ 0.05 200.28
FC-KAN (DoG+BS) 99.54 ± 0.13 89.99 ± 0.09 89.93 ± 0.08 369.2
FC-KAN (DoG+RBF) 99.68 ± 0.08 89.93 ± 0.07 89.87 ± 0.07 339.51
FC-KAN (DoG+BASE) 98.24 ± 0.55 89.81 ± 0.07 89.74 ± 0.07 326.38
FC-KAN (BS+RBF) 98.58 ± 0.43 89.72 ± 0.07 89.67 ± 0.06 327.49
FC-KAN (BS+BASE) 98.81 ± 0.35 89.75 ± 0.06 89.70 ± 0.06 327.31
FC-KAN (RBF+BASE) 98.97 ± 0.30 89.74 ± 0.05 89.69 ± 0.05 318.95
Train. Acc = Training Accuracy, Val. Acc. = Validation Accuracy
BASE = linear transformations, BS = B-splines, DoG = Difference of Gaussians, RBF = Radial Basis Functions
In general, FC-KAN models outperformed others on MNIST and Fashion-MNIST but require more training time due to the quadratic function representation for output combination, as shown in Table 2. This trade-off between training time and model performance is considered reasonable. The best-performing model is FC-KAN (DoG+BS), which achieved a validation accuracy of 97.91% on MNIST and 89.99% on Fashion-MNIST. Although it has the lowest performance on Fashion-MNIST, the MLP model has the fastest training time and shows competitive performance on MNIST, even outperforming BSRBF-KAN, FastKAN, FasterKAN, and EfficientKAN on this dataset.
BSRBF-KAN, FC-KAN (DoG+BS), and FC-KAN (DoG+RBF) exhibit the best convergence on MNIST, while FC-KAN (DoG+RBF) performs the best on Fashion-MNIST, followed by FC-KAN (DoG+BS) and BSRBF-KAN. We observe that fast convergence is achieved in KAN models that incorporate function combinations rather than relying on single functions. This finding is important to consider when designing KANs with a focus on achieving rapid convergence.
4.3 Comparison of Combination Methods
While employing a quadratic function representation for the output of FC-KAN, we are also interested in exploring how different output combination methods affect model performance. In this experiment, we use FC-KAN (DoG+BS) with several output combination methods: sum, element-wise product, addition of sum and element-wise product, quadratic and cubic function representations, concatenation, and linear transformation of the concatenated output. Inspired by the work of Altarabichi [11] on DropKAN [46], we also include the maximum, minimum, and average outputs for comparison.
Table 3: The performance of FC-KAN (DoG+BS) using different output combination methods.
Dataset Combined Method Train. Acc. Val. Acc. F1 Time (seconds)
MNIST Sum 100.00 ± 0.00 97.61 ± 0.04 97.58 ± 0.04 247.12
Product 100.00 ± 0.00 97.59 ± 0.07 97.56 ± 0.07 247.5
Sum + Product 100.00 ± 0.00 97.73 ± 0.04 97.70 ± 0.04 244.95
Quadratic Function 100.00 ± 0.00 97.91 ± 0.05 97.88 ± 0.05 263.29
Cubic Function 100.00 ± 0.00 97.68 ± 0.05 97.65 ± 0.05 266.46
Concatenation 99.64 ± 0.07 97.20 ± 0.02 97.16 ± 0.02 250.03
Linear Concatenation 100.00 ± 0.00 97.60 ± 0.04 97.56 ± 0.05 253.88
Minimum 99.92 ± 0.04 97.23 ± 0.04 97.20 ± 0.04 253.21
Maximum 99.97 ± 0.01 97.30 ± 0.05 97.26 ± 0.05 249.56
Average 100.00 ± 0.00 97.44 ± 0.02 97.40 ± 0.03 255.4
Fashion-MNIST Sum 99.39 ± 0.03 89.56 ± 0.07 89.55 ± 0.09 346.21
Product 99.50 ± 0.05 89.95 ± 0.08 89.90 ± 0.08 345.85
Sum + Product 99.56 ± 0.05 89.89 ± 0.13 89.84 ± 0.13 349.4
Quadratic Function 99.54 ± 0.13 89.99 ± 0.09 89.93 ± 0.08 369.2
Cubic Function 99.40 ± 0.05 89.69 ± 0.10 89.67 ± 0.09 367.83
Concatenation 95.27 ± 0.05 89.09 ± 0.04 89.01 ± 0.04 345.35
Linear Concatenation 99.53 ± 0.03 89.40 ± 0.06 89.37 ± 0.07 358.16
Minimum 98.13 ± 0.27 89.37 ± 0.06 89.33 ± 0.06 351.53
Maximum 97.47 ± 0.68 89.34 ± 0.06 89.28 ± 0.06 353.33
Average 99.09 ± 0.03 89.54 ± 0.06 89.51 ± 0.06 354.74
Train. Acc = Training Accuracy, Val. Acc. = Validation Accuracy
From the results in Table 3, the quadratic function best represents the combination output and outperforms other combinations, although its models require more training time. It is clear that the outputs combined by element-wise operations consistently outperform other methods, demonstrating superior accuracy. This indicates that element-wise combinations are more effective in capturing and integrating relevant features from the data, leading to better performance. Meanwhile, the output combination by concatenation shows the worst results. The linear transformation applied to the concatenated output outperforms concatenation alone, but still achieves only average performance. Similarly, the maximum, minimum, and average outputs do not deliver superior results.
In MNIST, besides the quadratic functions, the addition of sum and element-wise product demonstrates very competitive performance while requiring the least training time. Except for concatenation, maximum, and minimum, all other combinations can easily achieve 100% training accuracy. In Fashion-MNIST, the element-wise product combination is only surpassed by the quadratic function, but the plus point is it takes 6.3% less training time. The addition of sum and element-wise product achieves the best training accuracy, followed by the quadratic function and the linear transformation of concatenation.
Contrary to our expectations, the cubic function representation only achieves moderate performance. However, it takes the longest training time on MNIST and ranks just behind the quadratic function in terms of training time on Fashion-MNIST. Initially, we hypothesized that the cubic representation could capture more data features, but it appears that the excessive number of element-wise operations may hinder feature extraction, potentially leading to reduced performance. This experiment demonstrates that using higher-degree functions may not necessarily enhance model performance and can also increase computational complexity.
4.4 Model Performance with Limited Data
Rather than training FC-KAN models on the full data to determine the optimal configuration, we investigate whether using smaller portions of the data can yield comparable insights. This approach not only helps identify the best function combinations but also significantly reduces training time. Additionally, it allows for testing a greater number of configurations and provides early indications of model performance on the full training data. In this experiment, we evaluate 6 FC-KAN variants—FC-KAN (DoG+BS), FC-KAN (DoG+RBF), FC-KAN (DoG+BASE), FC-KAN (BS+RBF), FC-KAN (BS+BASE), and FC-KAN (RBF+BASE)—using 1%, 5%, and 10% of the data, with a quadratic function representation for combining function outputs. Each model is trained over 5 independent runs with the same configuration, and we calculate the average performance.
Figure 5: The validation accuracy values of the models across various data subsets.
Figure 5 illustrates the performance of the FC-KAN models across different training subsets and function combinations. In the MNIST dataset, the RBF+BASE combination demonstrates superior performance with 1%, 5%, and 10% of the data; however, the highest performance on the full training data is recorded for the DoG+BS combination. This observation underscores the challenges of predicting the optimal function combination when working with limited training data. In the Fashion-MNIST dataset, DoG+BS achieves the best performance with 1% of the data, while DoG+RBF excels with both 5% and 10%. These results suggest that DoG+BS and DoG+RBF may perform well with the full training dataset. Indeed, DoG+BS consistently outperforms other combinations in full training, with DoG+RBF serving as a strong competitor, as in Table 2. Overall, our findings indicate that accurately predicting the performance of FC variants on the complete training dataset based solely on their performance with smaller subsets is challenging. This variability may depend on the specific datasets and the portions of data used for training.
5 Limitation
Although FC-KAN is designed to utilize data combinations in low-dimensional layers, our experiments applied it only to the output layer, considered a low-dimensional layer in a network with the structure (784, 64, 10). As a result, the impact of these combinations on model performance in deeper network architectures with low-dimensional layers remains unclear. Another limitation is the number of parameters in the models. In the experiments, the MLP used the fewest parameters within the same network structure (784, 64, 10) compared to other models. We are also interested in how MLP would perform relative to KAN models if they have the same number of parameters. According to Yu et al. [62], MLP generally outperforms KAN models, except in tasks involving symbolic formula representation.
We also question whether the model’s performance would improve if data combinations were applied in all layers, rather than just low-dimensional layers, assuming that device memory constraints are not an issue in data multiplications. Finally, since FC-KAN has only been tested on two datasets, MNIST and Fashion-MNIST, more datasets should be used to properly evaluate its effectiveness. In short, these limitations can be addressed by designing network structures that integrate low-dimensional data and evaluating them across various problems or through additional experiments for greater clarity.
6 Conclusion
We introduced FC-KAN, which uses various popular mathematical functions to represent data features and combines their outputs using different methods, primarily through element-wise operations in low-dimensional layers, to address image classification problems. In the experiments, we designed FC-KAN to combine pairs of functions, such as B-splines, wavelets, and radial basis functions, using several output combinations on the MNIST and Fashion-MNIST datasets. We found that FC-KAN outperformed MLP and other KAN models with the same network structure, although it required more training time. Among the variants, FC-KAN (DoG+BS), which combines the Difference of Gaussians and B-splines with a quadratic function representation in the output, achieved the best results on both datasets. Our experiments also show that it is not easy to detect the function combination performance in the full training data based on small-scale experiments. FC-KAN demonstrates promising potential in using function combinations to design KANs and improve model performance. However, there are several aspects, such as exploring other functions and combinations to extract data features for further model enhancement, that we would like to investigate. These will be the focus of our future work.